# EditReward: A Human-Aligned Reward Model for Instruction-Guided Image Editing

### v1

@@ A.2 Details of EditReward-Data Construction
Quality Control. The reliability of our annotations is ensured through a multi-stage process. The process includes: (1) initial pilot studies to refine the annotation guidelines and rubric; (2) a formal training and calibration phase for all annotators to align their judgments; and (3) continuous random sampling and cross-checking of annotations during the formal labeling process to maintain a high inter-annotator agreement (IAA).
@@ A.3 Details of EditReward-Bench Construction
A.4 More details on Training Implementation
@@
Training Hyperparameters and Setup. The model was trained using the AdamW optimizer with the following configuration. All experiments were conducted using PyTorch 2.5.1 and Transformers 4.56.1.
  • Optimizer: AdamW ($\beta_{1}=0.9,\beta_{2}=0.95$)
  • Weight Decay: 0.1
  • Peak Learning Rate: $2\times 10^{-6}$
  • LR Schedule: Cosine decay with a warm-up ratio of 0.05.
  • Training Epochs: 2
  • Global Batch Size: 16 (Per-GPU batch size of 2 on 8 NVIDIA A100 GPUs).
  • Image Resolution: $448\times 448$ pixels (aspect ratio preserved).
  • Loss Function: Uncertainty-Aware Ranking Loss.
  • Aggregation Strategy: Balanced Average.
The 200K pairwise training samples were constructed from our expert annotations. For each source instruction, pairs of edited images were sampled, and their preference was determined by the aggregated point-wise scores. Our tie-disentanglement data augmentation strategy was applied to all qualifying tied pairs during training.

### v2

@@ A.2 Details of EditReward-Data Construction
Table 9: Inter-Annotator Agreement (IAA) Metrics for Expert Annotations. The robust Krippendorff Alpha ($\alpha$) values confirm the high reliability of our multi-dimensional expert scoring.
Dataset Fleiss’ Kappa (IF) Fleiss’ Kappa (VQ) Krippendorff Alpha (IF) Krippendorff Alpha (VQ)
EditReward-Data 0.4157 0.3203 0.6762 0.5720
EditReward-Bench 0.3962 0.3157 0.6623 0.6114
All data 0.3994 0.3111 0.6685 0.5972
Quality Control. The reliability of our annotations is ensured through a multi-stage process. The process includes: (1) initial pilot studies to refine the annotation guidelines and rubric; (2) a formal training and calibration phase for all annotators to align their judgments; and (3) continuous random sampling and cross-checking of annotations during the formal labeling process to maintain a high inter-annotator agreement (IAA). More representative examples from EditReward-Data are shown in the Appendix  A.10. The IAA results in Table 9 provide a critical quantitative assessment of our expert annotation quality (Fleiss, 1971). We highlight the values derived from Krippendorff’s Alpha ($\alpha$(Krippendorff, 2011), which is the most appropriate metric as it correctly models the ordinal nature of our 4-point Likert scale. The $\alpha$ scores of $\mathbf{0.668}$ for Instruction Following (IF) and $\mathbf{0.597}$ for Visual Quality (VQ) establish a strong, quantified baseline for human consistency. Crucially, the observed difference ($\text{IAA}_{\text{IF}}>\text{IAA}_{\text{VQ}}$) provides empirical validation for our core contribution: it confirms that the VQ dimension is inherently more subjective than IF. This validates our design choice to use a multi-dimensional rubric and a multi-head reward model, as a single holistic score would obscure this critical difference in human variance.
@@ A.3 Details of EditReward-Bench Construction
A.4 More details and implementation of training
@@ A.9 Input Template for Reward Model
A.10 Representative Results of EditReward-Data
The following examples provide additional qualitative illustrations of EditReward-Data. They highlight a broad spectrum of real editing behaviors, including appearance manipulation, object insertion/removal, style transfer and text change. Each example includes the source image, the edited result, and the associated annotations—such as instruction-following and visual quality. These samples complement the main paper by demonstrating the dataset’s diversity, annotation fidelity, and coverage across both everyday and challenging editing scenarios.
(a) Example 1 from EditReward-Data.
(b) Example 2 from EditReward-Data.
Figure 6: Representative examples from EditReward-Data, complementing Fig. 2.
(a) Example 3 from EditReward-Data.
(b) Example 4 from EditReward-Data.
Figure 7: Representative examples from EditReward-Data, complementing Fig. 2.
(a) Example 5 from EditReward-Data.
(b) Example 6 from EditReward-Data.
Figure 8: Representative examples from EditReward-Data, complementing Fig. 2.
A.11 Qualitative Examples of EditReward-Based Filtering
To provide additional intuition about the preferences learned by EditReward-Data ’s reward model, we visualize samples from the ShareGPT-4o-Image dataset (Chen et al., 2025a) that are either retained or filtered out after ranking with EditReward scores. The selected examples highlight the characteristic patterns captured by the reward model.
High-Quality Retained Data. Samples with high EditReward scores generally demonstrate accurate instruction following, clean and localized modifications, and visually coherent integration with the surrounding context. These images exhibit minimal artifacts and adhere closely to both the semantic intent and spatial constraints of the edit.
Low-Quality Filtered Data. Samples with low scores often contain undesirable visual artifacts, incorrect or incompletely executed edits, spatial misalignment, or hallucinated content. These failure patterns reflect typical challenges in image editing that violate instruction-following or degrade overall image quality.
Together, these qualitative examples illustrate the types of editing behaviors favored or penalized by EditReward, offering a clear and interpretable view of the model’s learned preferences during data filtering.
(a) Example List 1 of high-quality image editing samples.
(b) Example List 2 of high-quality image editing samples.
Figure 9: Examples of high-quality image editing samples that were retained after filtering based on EditReward scores.
(a) Example List 1 of low-quality image editing samples.
(b) Example List 2 of low-quality image editing samples.
Figure 10: Examples of low-quality image editing samples that were filtered out based on EditReward scores.
A.12 Qualitative Comparison: Before vs. After EditReward Filtering
To further illustrate the qualitative improvements enabled by EditReward-based data curation, we present side-by-side comparisons of image editing results produced by the same Step1X-Edit architecture trained on two datasets: (i) the original unfiltered dataset, and (ii) the EditReward-filtered high-quality subset.
For each example, we show the source image, the output from the model trained on unfiltered data (Before Filter), and the output from the model trained on EditReward-curated data (After Filter). These examples cover a range of editing types, including background replacement, object insertion/removal, style and material changes, and human-centric edits.
Qualitatively, the “After Filter” results exhibit more accurate instruction following, cleaner local modifications, fewer visual artifacts, and more coherent global integration. In contrast, models trained on unfiltered data tend to produce incomplete edits, spatial inconsistencies, or hallucinated structures. These comparisons highlight the alignment benefits of EditReward-guided data selection and validate its effectiveness in improving generation quality.
(a) Example 1: Comparison of edits produced before and after EditReward filtering.
(b) Example 2: Another comparison of Step1X-Edit results trained with and without EditReward filtering.
Figure 11: Qualitative comparison of Step1X-Edit models trained on unfiltered data (Before Filter) and EditReward-curated data (After Filter). Each example shows the source image, the output before filtering, and the output after filtering. Results demonstrate that training on EditReward-filtered data produces more accurate, stable, and faithful edits.
A.13 Failure Mode Analysis of EditReward
The following analysis moves beyond generic VLM weaknesses and focuses on specific, actionable biases exhibited by EditReward when its judgment significantly deviates from the human consensus. We found two persistent failure modes:
  1. 1.
    Visual Quality Bias (Color/Brightness): We observed that EditReward occasionally exhibits a perceptual bias toward vividness, likely inherited from its VLM pre-training. The model tends to conflate high overall visual quality with excessive brightness or high color saturation, resulting in inflated scores for edits that human judges deem over-processed or visually jarring.
  2. 2.
    Global Consistency Failure (Background Over-Editing): Despite our explicit training on the ”Exclusivity” criterion, the model sometimes gives high scores to edits where the background or unedited regions were substantially and unnecessarily altered. This suggests a specific weakness in balancing local edit success against the preservation mandate.
(a) Failure Mode 1: Brightness Bias. This image received a high score from EditReward (e.g., Score 3.8/4.0) due to its vivid colors, even though human experts rated it lower (e.g., Score 2.0) for being over-saturated and visually implausible. This illustrates the model’s tendency to reward excessive brightness.
(b) Failure Mode Examplex 2: Exclusivity/Background Failure. The edit (e.g., ”Change the object”) was successful locally, yet the model gave it a high reward (e.g., Score 3.5/4.0) despite the background being severely altered and distorted—a clear violation of the ”no unprompted changes” rule.
Figure 12: Qualitative Taxonomy of EditReward’s Reward Biases. These examples illustrate specific cases where EditReward’s high scores deviate from human consensus, demonstrating the model’s inherited bias toward high color vividness and its difficulty in penalizing subtle global inconsistencies.